# Tackling Interference Induced by Data Training Loops in A/B Tests: A Weighted Training Approach

### 03 Feb 2024 v4

@@ 1 Introduction
More specifically, let’s consider a user-side experiment testing two distinct ranking algorithms. In this scenario, we split the traffic in such a way that control users are subjected to control algorithms, and test users are subjected to test algorithms. Control and test algorithms generate data that may follow different distributions. These data sets are then combined and fed back into the ML models. This experimental procedure is represented in Figure 2.
@@ 2.1 Interference in Experiments
In addition to the papers mentioned above, interference has also been studied in other specialized settings. For instance, Chawla et al. (2016); Basse et al. (2016) and Liao and Kroer (2023) focus on experimental design in auctions. Han et al. (2023) employ roll-outs, a technique commonly implemented in experiments, to detect interference. Additionally, Boyarsky et al. (2023) demonstrate that roll-outs can also help estimation under stronger assumptions.
@@ 4 A Weighted Training Approach
1:The probability of treatment assignment: $p$; a model class for the weight prediction: $\mathcal{G}=\{G_{\theta_{W}}:\mathbb{R}^{d}\rightarrow\{0,1\},{\theta_{W}}\in\Theta_{W}\}$; the machine learning model class: =$\mathcal{M}=\{M_{\theta}:\mathcal{X}\rightarrow\mathcal{Y},\theta\in\Theta\}$; loss functions: $\ell(M(X),Y)$ (could be $m$-dimensional).
@@ 5.1 Simulation Setups
$\mathrm{FR}_{\mathrm{short}}$ $=$ $\beta_{\mathrm{FR,short}}^{\top}X-2.5,$
$\mathrm{FR}_{\mathrm{long}}$ $=$ $\beta_{\mathrm{FR,long}}^{\top}X-2.5,$
@@
$\mathrm{SD}_{\mathrm{long}}$ $\sim$ $\exp\left(\beta_{\mathrm{SD,long}}^{\top}X\right)$
where $\exp\left(\cdot\right)$ means an exponential distribution and
@@ 5.2 A/B Tests
We first examine the comparison between the control parameter $\alpha_{C}=10$ and the treatment parameter $\alpha_{T}=9$ with a treatment assignment probability of $p=1/2$. The results are depicted in Figure 4.
@@
Additionally, we provide information on the bias and standard errors of treatment effect estimators obtained using various methods in Table 1. In each metric, the first column represents the bias in comparison to the true global treatment effect (GTE). The second column displays the standard deviation calculated from the results of the 100 A/B tests. Lastly, the third column showcases the standard error estimates obtained through two-sample t-tests in a single A/B test i.e.,
@@
In Table 2, we have calculated the experimentation costs. For treatment users, we computed the average treatment values based on the treatment linear fusion formula, while for control users, we averaged the control values based on the control linear fusion formula. It’s apparent that our approach is only slightly worse than the global treatment/control regime, and the data splitting method incurs the lowest costs, indicating that it results in higher experimental expenses.

### 05 Apr 2024 v5

@@ 1 Introduction
More specifically, let’s consider a user-side experiment testing two distinct ranking algorithms. In this scenario, we split the traffic in such a way that control users are subjected to control algorithms, and treatment users are subjected to treatment algorithms. Control and treatment algorithms generate data that may follow different distributions. These data sets are then combined and fed back into the ML models. This experimental procedure is represented in Figure 2.
@@ 2.1 Interference in Experiments
In addition to the papers mentioned above, interference has also been studied in other specialized settings. For instance, Chawla et al. (2016); Basse et al. (2016); Liao and Kroer (2023), and Liao et al. (2024) focus on experimental design in auctions. Han et al. (2023) employ roll-outs, a technique commonly implemented in experiments, to detect interference. Additionally, Boyarsky et al. (2023) demonstrate that roll-outs can also help estimation under stronger assumptions.
@@ 4 A Weighted Training Approach
1:The probability of treatment assignment: $p$; a model class for the weight prediction: $\mathcal{G}=\{G_{\theta_{W}}:\mathbb{R}^{d}\rightarrow[0,1],{\theta_{W}}\in\Theta_{W}\}$; the machine learning model class: =$\mathcal{M}=\{M_{\theta}:\mathcal{X}\rightarrow\mathcal{Y},\theta\in\Theta\}$; loss functions: $\ell(M(X),Y)$ (could be $m$-dimensional).
@@ 5.1 Simulation Setups
$\mathrm{FR}_{\mathrm{short}}$ $=$ $\text{Sigmoid}(\beta_{\mathrm{FR,short}}^{\top}X-2.5),$
$\mathrm{FR}_{\mathrm{long}}$ $=$ $\text{Sigmoid}(\beta_{\mathrm{FR,long}}^{\top}X-2.5),$
@@
$\mathrm{SD}_{\mathrm{long}}$ $\sim$ $\exp\left(\beta_{\mathrm{SD,long}}^{\top}X\right),$
where $\mathrm{Sigmoid}$ means the sigmoid function and $\exp\left(\cdot\right)$ means an exponential distribution and
@@ 5.2 A/B Tests
We first examine the comparison between the control parameter $\alpha_{C}=10$ and the treatment parameter $\alpha_{T}=9$ with a treatment assignment probability of $p=1/2$. The logloss of our weighting model is about 0.96 (base 2). This value indicates a performance slightly above that of a purely random guess, which would yield a logloss of 1. Despite the marginal improvement over randomness in terms of logloss, it’s crucial to highlight that this translates into significant gains in the accuracy of treatment effect estimation, as we shall see in Figure 4 and Table 1.
@@
We further provide information on the bias and standard errors of treatment effect estimators obtained using various methods in Table 1. In each metric, the first column represents the bias in comparison to the true global treatment effect (GTE). The second column displays the standard deviation calculated from the results of the 100 A/B tests. Lastly, the third column showcases the standard error estimates obtained through two-sample t-tests in a single A/B test i.e.,
@@
In Table 2, we have calculated the experimentation costs. For treatment users, we computed the average treatment values based on the treatment linear fusion formula, i.e.,
$\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\alpha_{T}\text{Finish}_{i}+\text{StayDuration}_{i},$
where $N_{T}$ represents the number of users in the treatment group and $\text{Finish}_{i}$ and $\text{StayDuration}_{i}$ indicate whether a user finished watching a video and their duration of stay, respectively. While for control users, we averaged the control values based on the control linear fusion formula:
$\frac{1}{N_{C}}\sum_{i=1}^{N_{C}}\alpha_{C}\text{Finish}_{i}+\text{StayDuration}_{i}.$
It’s apparent that our approach is only slightly worse than the global treatment/control regime, and the data splitting method incurs the lowest costs, indicating that it results in higher experimental expenses.